# MoM: Linear Sequence Modeling with Mixture-of-Memories

### v2

@@ Abstract
Linear sequence modeling methods, such as linear attention, state space modeling, and linear RNNs, offer significant efficiency improvements by reducing the complexity of training and inference. However, these methods typically compress the entire input sequence into a single fixed-size memory state, which leads to suboptimal performance on recall-intensive downstream tasks. Drawing inspiration from neuroscience, particularly the brain’s ability to maintain robust long-term memory while mitigating "memory interference", we introduce a novel architecture called Mixture-of-Memories (MoM). MoM utilizes multiple independent memory states, with a router network directing input tokens to specific memory states. This approach greatly enhances the overall memory capacity while minimizing memory interference. As a result, MoM performs exceptionally well on recall-intensive tasks, surpassing existing linear sequence modeling techniques. Despite incorporating multiple memory states, the computation of each memory state remains linear in complexity, allowing MoM to retain the linear-complexity advantage during training, while constant-complexity during inference. Our experimental results show that MoM significantly outperforms current linear sequence models on downstream language tasks, particularly recall-intensive tasks, and even achieves performance comparable to Transformer models.
The code is released at https://github.com/OpenSparseLLMs/MoM and is also released as a part of https://github.com/OpenSparseLLMs/Linear-MoE.
MoM: Linear Sequence Modeling with Mixture-of-Memories
Jusen Du1,2, Weigao Sun1🖂, Disen Lan1,3, Jiaxi Hu4, Yu Cheng5🖂 1Shanghai AI Laboratory, 2Nanjing University, 3South China University of Technology, 4The Hong Kong University of Science and Technology (Guangzhou), 5The Chinese University of Hong Kong
"The enemy of memory isn’t time; it’s other memories."
— David Eagleman
† interns at Shanghai AI Laboratory; 🖂  corresponding authors (sunweigao@outlook.com, chengyu@cse.cuhk.edu.hk). Weigao Sun is the project lead.
@@ 1 Introduction
Attention mechanisms have made significant contributions to the field of artificial intelligence, advancing various modalities such as language, vision, audio, video, graphs, and even time series Achiam et al. (2023); Team (2023). The Transformer Vaswani (2017), known for its ability to capture long-range dependencies, has become a foundational architecture in this space. However, traditional Transformers encounter computational challenges due to their quadratic time complexity, $O(n^{2})$, with respect to sequence length $n$, making it difficult to scale to long sequences. To overcome this limitation, several linear sequence modeling methods have been proposed, including linear attention Katharopoulos et al. (2020); Qin et al. (2023a); Li et al. (2025), state space modeling Gu and Dao (2024); Dao and Gu (2024), and linear RNNs Peng et al. (2024); Qin et al. (2024d), which offer $O(n)$ training complexity and $O(1)$ inference complexity. These approaches often reduce the input sequence to a fixed-size hidden space, collapsing the information into a single "memory state". While these methods enhance efficiency, they face two main challenges: limited memory capacity and memory interference. When new information overwrites the single fixed-size memory state, previously stored representations may degrade, which negatively impacts its long-term memory performance on recall-intensive tasks.
We argue that the strong performance of Transformer models on recall-intensive tasks arises from their ability to avoid "memory interference" by maintaining independent key-value caches for each token, thus offering virtually unlimited memory capacity. In contrast, linear sequence modeling relies on extreme compression, consolidating all the input information into a single fixed-size memory state Katharopoulos et al. (2020); Dao and Gu (2024). This approach results in limited memory capacity and inherently leads to memory interference issues.
Interestingly, the human brain has developed mechanisms that enable large memory capacity while reducing memory interference. Neuroscience studies show that in the hippocampus, theta oscillations (4$\sim$8 Hz) and gamma oscillations (30$\sim$100 Hz) work together to support a neural coding mechanism for multi-item memory Buzsáki (2002); Lisman and Jensen (2013). Specifically, each theta cycle is subdivided into multiple gamma subcycles, and within each gamma subcycle, a distinct group of neurons is activated following the "E%-max" mechanism de Almeida et al. (2009). This sequential activation temporally separates different memory items, thus preventing interference.
@@
  • We present MoM, a neuroscience-inspired architecture that incorporates multiple independent memory states, significantly enhancing memory capacity and eliminating memory interference, while retaining the efficiency benefits of linear-time training and constant-memory inference.
  • Distinct with existing gating mechanisms, MoM is a new paradigm to reduce memory interference by separating the memory states. The overall design of MoM is compatible with existing linear sequence modeling methods, making it a straightforward and effective approach to boost task performance.
  • Through empirical evaluation, we show that MoM outperforms state-of-the-art linear sequence modeling techniques across a variety of language tasks, particularly on recall-intensive tasks. MoM even achieves performance on par with Transformer models, a feat that current linear sequence modeling methods struggle to match.
@@ 2 Preliminary
To reduce the time complexity of Transformer attention, various optimization techniques have been proposed. Linear Transformers Katharopoulos et al. (2020) replace the softmax attention mechanism with dot-product of feature maps $\phi(\cdot)$:
@@
where $\bm{q}_{t},\bm{k}_{t},\bm{v}_{t}\in\mathbb{R}^{d}$. While the presence of the denominator may lead to numerical instability Qin et al. (2024b) and the feature map can utilize an identity function, which we omit for simplicity. In perspective of memory, the formulation can also be written in a recurrent format:
$\bm{M}_{t}=\bm{M}_{t-1}+\bm{k}_{t}^{T}\bm{v}_{t},\quad\bm{o}_{t}=\bm{q}_{t}\bm{M}_{t},$ (2)
@@
Building on the foundational concepts of linear attention and memory perspective, some recent advancements have focused on optimizing memory structure, including gated updates Yang et al. (2023); Qin et al. (2024e, d) and memory capacity expansion Peng et al. (2024); Qin et al. (2024d).
@@ 3.1 Motivation
Linear sequence models compress the entire sequence data into a fixed-size memory state. Despite numerous efforts to minimize information loss—such as introducing gating mechanisms and employing more precise control over memory modifications Orvieto et al. (2023); De et al. (2024); Beck et al. (2024); Yang et al. (2023); Zhang et al. (2024)—some degradation in this compression process is inevitable. Expanding the memory capacity has been shown to mitigate this issue to some extent, with studies indicating that increasing memory capacity can enhance model performance Qin et al. (2024d); Peng et al. (2024).
@@ 3.2 MoM: Mixture-of-Memories
To address the challenge outlined above, we propose a novel approach inspired by biological mechanisms for encoding multi-item memory such as theta-gamma oscillations Lisman and Jensen (2013), and concepts from Mixture-of-Experts (MoE) Shazeer et al. (2017), where different experts handle specific tokens. In this approach, we leverage multiple memory states, each of which is selectively updated by different inputs. This increases the memory capacity and enables the model to retain diverse pieces of information by storing various types of inputs in separate memory states.
In our framework, the memory states function similarly to the experts in MoE. However, instead of relying on completely separate networks, these modules are individual RNN states embedded within a linear recurrent mechanism. This design allows for the isolation of memory updates while concurrently managing distinct types of information. It is important to note that MoM fundamentally differs from traditional MoE, as we will discuss in Appendix B. Figure 1 provides an overview of the MoM architecture. Below, we introduce the structure of the MoM layer and explain how this biologically-inspired architecture is implemented in the context of linear sequence modeling.
@@ 3.2.2 Linear Recurrent Memory Module
    $\bm{M}^{m}_{t}=\bm{M}^{m}_{t-1}+(\bm{k}^{m}_{t})^{T}\bm{v}^{m}_{t}\in\mathbb{R}^{d\times d},$ (6)
    The equation above represents the simplest form of memory update for clarity. Our approach is flexible and does not rely on a specific memory update mechanism. To enhance performance, we can incorporate mechanisms such as forget gates Sun et al. (2023):
    $\bm{M}^{m}_{t}=\gamma\bm{M}^{m}_{t-1}+(\bm{k}^{m}_{t})^{T}\bm{v}^{m}_{t}\in\mathbb{R}^{d\times d},$ (7)
    where $\gamma$ is a constant forget gate.
@@
Method Memory Update Rule
Linear Attn $\bm{M}_{t}=\bm{M}_{t-1}+\bm{k}_{t}^{T}\bm{v}_{t}$
Lightning $\bm{M}_{t}=\gamma\bm{M}_{t-1}+\bm{k}_{t}^{T}\bm{v}_{t}$
RetNet $\bm{M}_{t}=\gamma\bm{M}_{t-1}+\bm{k}_{t}^{T}\bm{v}_{t}$
GLA $\bm{M}_{t}=(\bm{a}_{t}^{T}\bm{1})\bm{M}_{t-1}+\bm{k}_{t}^{T}\bm{v}_{t}$
DeltaNet $\bm{M}_{t}=(\bm{I}-\bm{k}_{t}^{T}\bm{k}_{t})\bm{M}_{t-1}+b_{t}\bm{k}_{t}^{T}\bm{v}_{t}$
G-DeltaNet $\bm{M}_{t}=a_{t}(\bm{I}-\bm{k}_{t}^{T}\bm{k}_{t})\bm{M}_{t-1}+b_{t}\bm{k}_{t}^{T}\bm{v}_{t}$
TTT $\bm{M}_{t}=\bm{M}_{t-1}+b_{t}\nabla l(\bm{M}_{t-1};\bm{k}_{t},\bm{v}_{t})$
Titans* $\bm{M}_{t}=a_{t}\bm{M}_{t-1}+b_{t}\nabla l(\bm{M}_{t-1};\bm{k}_{t},\bm{v}_{t})$
Mamba2 $\bm{M}_{t}=a_{t}\bm{M}_{t-1}+b_{t}\bm{k}_{t}^{T}\bm{v}_{t}$
HGRN2 $\bm{M}_{t}=(\bm{a}_{t}^{T}\bm{1})\bm{M}_{t-1}+(1-\bm{a}_{t})^{T}\bm{v}_{t}$
RWKV6 $\bm{M}_{t}=a_{t}\bm{M}_{t-1}+\bm{k}_{t}^{T}\bm{v}_{t}$
RWKV7 $\bm{M}_{t}=a_{t}\bm{M}_{t-1}+b_{t}\nabla l(\bm{M}_{t-1};\bm{k}_{t},\bm{v}_{t})$
    For Titans, we have omitted details like momentum and norm optimization during gradient descent to keep the expression simplified.
Table 1: Memory Update Rules. We demonstrate that several current linear sequence models can be viewed as recurrent models in terms of memory updates, where $a_{t},b_{t}\in(0,1)$ are data-dependent scaler, $\bm{a}_{t}$ is data-dependent vector, and $\gamma$ is a data-independent constant.
@@
where $\bm{M}_{m}$ is one activated memory and $g_{t}^{(m)}$ is the importance score of $\bm{M}_{m}$.
@@
Finally, the output of the MoM layer is computed by applying an activation function, normalization, and a linear transformation:
$\bm{o}_{t}=\mathrm{RMSNorm}(\mathrm{Swish}(\bm{o}_{t}))\bm{W}_{o}\in\mathbb{R}^{d},$ (10)
@@
Throughout the recurrent process, only a subset of memory states is activated and updated at each time step, while memory states that are not routed remain inactive and unchanged. When the input passes through the key-value projection layer, it generates multiple sets of keys and values that are fed into different memory modules. This design enables the model to maintain multiple memory states, each preserving distinct pieces of information. By aggregating the activated memories into a comprehensive mixed memory by weighted summation, the query can effectively retrieve information from this mixed memory, which results the "attention output" followed by other layers.

### v3

@@ Abstract
Linear sequence modeling methods, such as linear attention, state space modeling, and linear RNNs, offer significant efficiency improvements by reducing the complexity of training and inference. However, these methods typically compress the entire input sequence into a single fixed-size memory state, which leads to suboptimal performance on recall-intensive tasks. To address this limitation, we introduce a novel architecture called Mixture-of-Memories (MoM). MoM utilizes multiple independent memory states, with a router network directing input tokens to specific memory states. This approach greatly enhances the overall memory capacity while minimizing memory interference. MoM serves as a general framework that can be seamlessly combined with diverse memory update mechanisms across linear models. As a result, MoM performs exceptionally well on recall-intensive tasks, surpassing existing linear sequence modeling techniques. Despite incorporating multiple memory states, the computation of each memory state remains linear in complexity, allowing MoM to retain the linear-complexity advantage during training, while constant-complexity during inference. Our experimental results show that MoM outperforms current linear sequence models on downstream language tasks, particularly recall-intensive tasks, and even achieves performance comparable to Transformer models. The code is released at https://github.com/OpenSparseLLMs/MoM and is also released as a part of https://github.com/OpenSparseLLMs/Linear-MoE.
* Interns at Shanghai AI Laboratory; 🖂 Corresponding Authors: Weigao Sun (sunweigao@outlook.com) and Yu Cheng (chengyu@cse.cuhk.edu.hk); † Project Lead.
”The enemy of memory isn’t time; it’s other memories.”
– David Eagleman
@@ 1 Introduction
Attention mechanisms have made significant contributions to the field of artificial intelligence, advancing various modalities such as language, vision, audio, video, graphs, and even time series (Achiam et al., 2023; Team, 2023). The Transformer (Vaswani, 2017), known for its ability to capture long-range dependencies, has become a foundational architecture in this space. However, traditional Transformers encounter computational challenges due to their quadratic time complexity, $O(n^{2})$, with respect to sequence length $n$, making it difficult to scale to long sequences. To overcome this limitation, several linear sequence modeling methods have been proposed, including linear attention (Katharopoulos et al., 2020; Qin et al., 2023a; Li et al., 2025), state space modeling (Gu & Dao, 2024; Dao & Gu, 2024), and linear RNNs (Peng et al., 2024; Qin et al., 2024d), which offer $O(n)$ training complexity and $O(1)$ inference complexity. These approaches often reduce the input sequence to a fixed-size hidden space, collapsing the information into a single “memory state”. While these methods enhance efficiency, they face two main challenges: limited memory capacity and memory interference. When new information overwrites the single fixed-size memory state, previously stored representations may degrade, which negatively impacts its long-term memory performance on recall-intensive tasks.
We argue that the strong performance of Transformer models on recall-intensive tasks arises from their ability to avoid memory interference by maintaining independent key-value caches for each token, thus offering virtually unlimited memory capacity. In contrast, linear sequence modeling relies on extreme compression, consolidating all the input information into a single fixed-size memory state (Katharopoulos et al., 2020; Dao & Gu, 2024). This approach results in limited memory capacity and inherently leads to memory interference issues.
[…]
@@
  • We present MoM, an architecture that incorporates multiple independent memory states, significantly enhancing memory capacity and eliminating memory interference, while retaining the efficiency benefits of linear-time training and constant-memory inference.
  • Distinct with existing gating mechanisms, MoM is a new paradigm to reduce memory interference by separating the memory states. The overall design is broadly compatible with diverse linear sequence modeling methods, making it a straightforward and effective approach to boost task performance.
  • Through empirical evaluation, we show that MoM outperforms strong linear sequence modeling baselines across a variety of language tasks, particularly on recall-intensive tasks. MoM even achieves performance on par with Transformer models, a feat that current linear sequence modeling methods struggle to match.
@@ 2 Preliminary
To reduce the time complexity of Transformer attention, various optimization techniques have been proposed. Linear Transformers (Katharopoulos et al., 2020) replace the softmax attention mechanism with dot-product of feature maps $\phi(\cdot)$:
@@
where $\bm{q}_{t},\bm{k}_{t},\bm{v}_{t}\in\mathbb{R}^{d}$. While the presence of the denominator may lead to numerical instability (Qin et al., 2024b) and the feature map can utilize an identity function, which we omit for simplicity. In perspective of memory, the formulation can also be written in a recurrent format:
$\bm{M}_{t}=\bm{M}_{t-1}+\bm{k}_{t}^{T}\bm{v}_{t},\quad\bm{o}_{t}=\bm{q}_{t}\bm{M}_{t}.$ (2)
@@
Building on the foundational concepts of linear attention and memory perspective, some recent advancements have focused on optimizing memory structure, including gated updates (Yang et al., 2023; Qin et al., 2024e; d) and memory capacity expansion (Peng et al., 2024; Qin et al., 2024d).
@@ 3.1 Motivation
Linear sequence models compress the entire sequence data into a fixed-size memory state. Despite numerous efforts to minimize information loss, such as introducing gating mechanisms and employing more precise control over memory modifications (Orvieto et al., 2023; De et al., 2024; Beck et al., 2024; Yang et al., 2023; Zhang et al., 2024), some degradation in this compression process is inevitable. Expanding the memory capacity has been shown to mitigate this issue to some extent, with studies indicating that increasing memory capacity can enhance model performance (Qin et al., 2024d; Peng et al., 2024).
@@ 3.2 MoM: Mixture-of-Memories
To address the challenge outlined above, we propose a novel approach for encoding multi-item memory such as theta-gamma oscillations (Lisman & Jensen, 2013), and concepts from Mixture-of-Experts (MoE) (Shazeer et al., 2017), where different experts handle specific tokens. In this approach, we leverage multiple memory states, each of which is selectively updated by different inputs. This increases the memory capacity and enables the model to retain diverse pieces of information by storing various types of inputs in separate memory states.
In our framework, the memory states function similarly to the experts in MoE. However, instead of relying on completely separate networks, these modules are individual RNN states embedded within a linear recurrent mechanism. This design allows for the isolation of memory updates while concurrently managing distinct types of information. It is important to note that MoM fundamentally differs from traditional MoE, as we will discuss in Appendix B. Figure 1 provides an overview of the MoM architecture. Below, we introduce the structure of the MoM layer and explain how this multi-memory architecture is implemented in the context of linear sequence modeling.
@@ 3.2.2 Linear Recurrent Memory Module
Table 1: Memory Update Rules. We demonstrate that several linear sequence models can be viewed as recurrent models in terms of memory updates, where $a_{t},b_{t}\in(0,1)$ are data-dependent scaler, $\bm{a}_{t}$ is data-dependent vector, and $\gamma$ is a data-independent constant.
Method Memory Update Rule Linear Attn $\bm{M}_{t}=\bm{M}_{t-1}+\bm{k}_{t}^{T}\bm{v}_{t}$ RetNet $\bm{M}_{t}=\gamma\bm{M}_{t-1}+\bm{k}_{t}^{T}\bm{v}_{t}$ GLA $\bm{M}_{t}=(\bm{a}_{t}^{T}\bm{1})\bm{M}_{t-1}+\bm{k}_{t}^{T}\bm{v}_{t}$ DeltaNet $\bm{M}_{t}=(\bm{I}-\bm{k}_{t}^{T}\bm{k}_{t})\bm{M}_{t-1}+b_{t}\bm{k}_{t}^{T}\bm{v}_{t}$ G-DeltaNet $\bm{M}_{t}=a_{t}(\bm{I}-\bm{k}_{t}^{T}\bm{k}_{t})\bm{M}_{t-1}+b_{t}\bm{k}_{t}^{T}\bm{v}_{t}$ TTT $\bm{M}_{t}=\bm{M}_{t-1}+b_{t}\nabla l(\bm{M}_{t-1};\bm{k}_{t},\bm{v}_{t})$ Titans $\bm{M}_{t}=a_{t}\bm{M}_{t-1}+b_{t}\nabla_{M}l(\bm{M}_{t-1};\bm{k}_{t},\bm{v}_{t})$ Mamba2 $\bm{M}_{t}=a_{t}\bm{M}_{t-1}+b_{t}\bm{k}_{t}^{T}\bm{v}_{t}$ HGRN2 $\bm{M}_{t}=(\bm{a}_{t}^{T}\bm{1})\bm{M}_{t-1}+(1-\bm{a}_{t})^{T}\bm{v}_{t}$ RWKV6 $\bm{M}_{t}=a_{t}\bm{M}_{t-1}+\bm{k}_{t}^{T}\bm{v}_{t}$ RWKV7 $\bm{M}_{t}=(\bm{a}_{t}^{T}\bm{1})\bm{M}_{t-1}+b_{t}\nabla l(\bm{M}_{t-1};\bm{k}_{t},\bm{v}_{t})$
@@
    $\bm{M}^{m}_{t}=\bm{M}^{m}_{t-1}+(\bm{k}^{m}_{t})^{T}\bm{v}^{m}_{t}\in\mathbb{R}^{d\times d}.$ (6)
    The equation above represents the simplest form of memory update for clarity. Our approach is flexible and does not rely on a specific memory update mechanism. To enhance performance, we can incorporate mechanisms such as forget gates (Sun et al., 2023).
@@
where $\bm{M}^{m}_{t}$ is one activated memory and $g_{t}^{(m)}$ is the importance score of $\bm{M}^{m}_{t}$.
@@
Finally, the output of the MoM layer is computed by applying an activation function, normalization, and a linear transformation.
@@
Throughout the recurrent process, only a subset of memory states is activated and updated at each time step, while memory states that are not routed remain inactive and unchanged. When the input passes through the key-value projection layer, it generates multiple sets of keys and values that are fed into different memory modules. This design enables the model to maintain multiple memory states, each preserving distinct pieces of information. By aggregating the activated memories into a comprehensive mixed memory by weighted summation, the query can effectively retrieve information from this mixed memory, and generate attention output followed by other layers.
Figure 2: Hardware-efficient Implementation of MoM. Tokens sharing the same color are routed to the same memory. ① Tokens are first split into groups according to memory routing results, ② then concatenated into a varlen input sequence, ③ processed by the Triton kernel, ④ the outputs are returned, ⑤ split back into their respective memories, and ⑥ finally restored to the original sequence order. For clarity, the illustration shows the top-$1$ routing case, and the qkvqkv projection is omitted.
3.3 Hardware-Efficient Implementation
In the implementation of MoM, mixing memories before query multiplication is equivalent to multiplying each memory by the query and then mixing the results, allowing us to reuse efficient Triton-based operators from prior linear sequence models. We first reorder the sequence tokens according to the routing results so that they follow the memory layout. The reordered tokens are then concatenated with varlen for operator computation, after which the results are aggregated via weighted summation. In this way, MoM’s computation can be effectively reduced to varlen operations, enabling efficient execution. We elaborate on this process below.
Given input tokens $\bm{x}_{b,t}\in\mathbb{R}^{d}$ for batch $b\in\{1,\dots,B\}$ and time step $t\in\{1,\dots,T\}$, each token is routed to one or more memories $m\in\{1,\dots,M\}$ with routing weights $\alpha_{b,t,m}\geq 0$ satisfying $\sum_{m=1}^{M}\alpha_{b,t,m}=1$.
For each $(b,m)$, define the ordered index set
$\mathcal{I}_{b,m}=\bigl(t_{b,m}(1),\dots,t_{b,m}(L_{b,m})\bigr),$
where $t_{b,m}(j)$ is the original sequence index of the $j$-th token assigned to memory $m$, and $L_{b,m}=|\mathcal{I}_{b,m}|$. We index buckets lexicographically by p=(b1)M+mp=(b-1)M+m and define cumulative boundaries
$s_{0}=0,\qquad s_{p}=\sum_{q=1}^{p}L_{q}\quad(p=1,\dots,BM).$
The flattened sequence $\bm{\tilde{X}}$ is obtained by
$\bm{\tilde{x}}_{\,s_{p-1}+j}=\bm{x}_{\,b,\,t_{b,m}(j)},\qquad j=1,\dots,L_{b,m},$
with varlen representation $(\bm{\tilde{X}},\bm{s})$, where $\bm{s}=(s_{0},\dots,s_{BM})$.
For each bucket p=(b1)M+mp=(b-1)M+m, queries share a projection matrix $\bm{W}_{Q}$, while keys and values use memory-specific projections $\bm{W}_{K}^{(m)},\bm{W}_{V}^{(m)}$:
$\bm{\tilde{q}}_{u}=\bm{W}_{Q}\bm{\tilde{x}}_{u},\quad\bm{\tilde{k}}_{u}=\bm{W}_{K}^{(m)}\bm{\tilde{x}}_{u},\quad\bm{\tilde{v}}_{u}=\bm{W}_{V}^{(m)}\bm{\tilde{x}}_{u},\quad u\in\{s_{p-1}+1,\dots,s_{p}\}.$
A memory-specific kernel $\mathcal{F}_{m}$ with parameters $\bm{\theta}^{(m)}$ is applied independently to each segment:
$\bm{o}_{\,s_{p-1}+1:s_{p}}=\mathcal{F}_{m}\!\bigl(\bm{\tilde{q}}_{s_{p-1}+1:s_{p}},\,\bm{\tilde{k}}_{s_{p-1}+1:s_{p}},\,\bm{\tilde{v}}_{s_{p-1}+1:s_{p}};\,\bm{\theta}^{(m)}\bigr).$
Mapping outputs back to the original sequence, the $j$-th token in $\mathcal{I}_{b,m}$ has per-memory output
$\bm{\hat{o}}_{b,t_{b,m}(j),m}=\bm{o}_{\,s_{p-1}+j}.$
Finally, token-level representations are reconstructed by weighted summation:
$\bm{y}_{b,t}=\sum_{m=1}^{M}\alpha_{b,t,m}\,\bm{\hat{o}}_{b,t,m}.$
4 Experiments
4.1 Experimental Setups
Models.  In our experiments, we employ the Gated DeltaNet (Yang et al., 2024) as the memory update mechanism in MoM. The model is configured with four memory states, two of which are activated at each time step, along with a shared memory.
Baselines.  We evaluate MoM against several linear recurrent models and Transformers, including RetNet (Sun et al., 2023), GLA (Yang et al., 2023), Gated DeltaNet (Yang et al., 2024), and Transformer++ (Touvron et al., 2023), which incorporates Rotary Position Embeddings (Su et al., 2024) and GLU (Shazeer, 2020) into the Transformer architecture. To ensure a fair comparison, we train all baseline models from scratch using the exact same number of tokens.
Training.  We follow the training procedure described by Yang et al. (2023), utilizing the SlimPajama dataset (Soboleva et al., 2023) sampled with 100B tokens and tokenized using the Mistral tokenizer (Jiang et al., 2023). We train models from scratch with parameter sizes of 380M and 1.3B, respectively. For the 380M models, we train on 15B tokens with a batch size of 0.5M tokens. More detailed training configuration is provided in Appendix C. We utilized publicly available pretrained weights from Zhang et al. (2024) with exactly same configuration  [^1] Models marked with an asterisk use open-source pretrained weights with identical training configurations..
Parameter Explanation. We report model sizes using the common shorthand, where “380M” denotes a configuration with 24 layers and hidden size 1024, and “1.3B” denotes 24 layers with hidden size 2048. The main goal of MoM is to expand the memory capacity of linear sequence models through sparse activation. To this end, we apply sparse activation only to the key and value projections, which results in a small increase in activated parameters that is well justified by the performance gains. A detailed discussion on fairness is provided in Appendix G.
4.2 Main Results